Ada and Stan are peeling a pile of potatoes. Working alone it would take Ada 1 hour to peel all the potatoes and it would take Stan 1 1/3 hours to peel all the potatoes. How long will it take to peel the potatoes if Ada and Stan work together? What portion of the pile does each peel?
Answer by lwsshak3(11628)

Ada and Stan are peeling a pile of potatoes. Working alone it would take Ada 1 hour to peel all the potatoes and it would take Stan 1 1/3 hours to peel all the potatoes. How long will it take to peel the potatoes if Ada and Stan work together? What portion of the pile does each peel
**
1/1=Ada's work rate
1/(4/3)=Stan's work rate
1/x=work rate when Ada and Stan work together
..
sum of individual work rates=work rate when working together
1+3/4=1/x
x+3x/4=1
4x+3x=4
7x=4
x=4/7 hrs to complete the job working together
1*4/7=4/7=Ada's portion
3/4*4/7=12/28=3/7=Stan's portion